1. Which of the following is a correct statement
A. 0.03>30% B. 0.03=3/10 C. 0.30<3/10 D. 30%=3/10

Answer by stanbon(48546)

Which of the following is a correct statement
A. 0.03>30%
false
---------------------
B. 0.03=3/10
false
---------------------
C. 0.30<3/10
false
----------------
D. 30%=3/10
30/100 = 3/10
true
==============
Cheers,
Stan H.